OpenStudy (anonymous):

12.8 is 32% of what number.show. The equation

OpenStudy (cornzyblack):

(32/100) * a= 12.8 0.32a= 12.8 a=12.8/0.32 a= 40

OpenStudy (anonymous):

12.8 over x = 32 over 100. cross multiply and get 1280=32x. divide by 32 and get 40. Check…0.32 times 40 = 12.8